# Arrival Time Prediction for Autonomous Shuttle Services in the Real World: Evidence from Five Cities

### v1

@@ V-A Data pre-processing
Finally, we enrich our data set with weather data at an hourly resolution and perform sine and cosine encodings of time-of-day and day-of-week features. For a correct model evaluation, we hold out a part of the data set depending on the date to avoid data leakage. In Linköping, Tampere, Madrid, and Graz, our test set selects one month of data and one week in Les Mureaux.
@@
The data pre-processing step is of utmost importance to make an accurate distinction between dwell time and running time. Since our data originates from pilot tests, careful data cleaning is essential to obtain a data set that reflects regular operation as accurately as possible.
VI Results
In this section, we perform a series of experiments on the data sets presented in the previous section. First, we investigate whether it is beneficial to create historical lags based on all vehicles or per vehicle. Then we compare the different model’s performance in dwell time and running time predictions. Lastly, we aggregate the predictions from the models to calculate AT predictions along observed journeys from the test data to evaluate the aggregation error.
As baselines for our prediction models, we compare the performance of the Lag model, which predicts the most recent observation, and the Mean regressor, also often referred to as a historical average model, which predicts the mean of each segment. The hyperparameters and training scripts for all evaluated models are available.  [^1] https://github.com/carolinssc/Arrival-time-prediction-autonomous-shuttles As performance metrics, we report the Root Mean Square Error (RMSE) as well as the Mean Absolute Error (MAE):
$\text{RMSE}(y,\hat{y})$ $=\sqrt{\frac{\sum_{i=0}^{N-1}(y_{i}-\hat{y}_{i})^{2}}{N}}$ (4)
$\text{MAE}(y,\hat{y})$ $=\frac{\sum_{i=0}^{N-1}|y_{i}-\hat{y}_{i}|}{N}$ (5)
where $y$ is the true labels and $\hat{y}$ is the predictions for the test set with $N$ observations.
Prior to assessing the final prediction results, we need to determine whether lags are taken from the most recent observation of the individual vehicle or the whole fleet. Hence, in Table II, we compare the predictive performance of the Lag Model and XGBoost on running time depending on the lags. Pilot testing occasionally results in sparse data, making recent lags from the whole fleet potentially advantageous. If, on the other hand, the fleet is a mix of shuttles from different manufacturers, observations from different vehicles may not be as informative. As seen in Table II, the results are not consistent across sites. Les Mureaux and Tampere deploy the same type of vehicles, while the remaining sites test different manufacturers. As a consequence, we can see a clear improvement in Linköping, with the highest number of vehicles deployed, when providing lags from the respective vehicle. For the other pilot sites, the difference is less striking. In Les Mureaux, the vehicles are identical, which leads to a better lag model for the lags based on the whole fleet. Yet, XGBoost performs better at the individual vehicle level, indicating subtle differences in the data of the vehicles. Based on this analysis, we create the lags per vehicle in Linköping, Tampere, and Les Mureaux and for the whole fleet in Madrid.
@@ VI Results
TABLE II: Results for Running Time Predictions Depending on Lags (Results are (RMSE/MAE)
@@
In Table III, we present the performance for dwell time predictions of the different models. Interestingly, there is no universal best model across all sites. In Linköping and Tampere, XGBoost outperforms in RMSE, but RF-GCN achieves the lowest MAE, while the difference in MAE is more significant than in RMSE. This implies that integrating a GNN with an upstream classifier could be the best strategy for these locations. The difference in performance between MLP and GCN, especially in Linköping, underlines the benefits of integrating spatial information, specifically for pilot sites with good data availability and longer, frequently-used routes. In Les Mureaux, the situation is reversed. The tree-based models outperform the other models by a large margin. We perform a more detailed analysis of this difference in the subsequent section. Given the very limited data availability in Graz, it is no surprise that our tested models cannot outperform the mean regressor, as there is not enough data for the advanced models to learn the spatial and temporal correlations. In Madrid, the GCN exhibits better performance than RF-GCN. We can attribute this to an observed degradation in the performance of the RF classifier in detecting zeros compared to the standalone GCN. This finding is intriguing, suggesting that spatial information in this pilot not only helps in accurate dwell time prediction when a shuttle stops but also in detecting stop skips.
TABLE III: Results for Dwell Time Prediction
Lag Mean LR RF XGBoost MLP GCN RF-GCN
MAE 5.41 8.51 5.85 3.30 3.46 $5.45\pm 0.11$ $5.23\pm 0.15$ $3.98\pm 0.13$
Graz RMSE 15.84 10.88 11.27 10.52 10.66 $12.71\pm 0.24$ $12.92\pm 0.17$ $11.21\pm 0.09$
In Table IV, we analyze if the results from the dwell time predictions translate to running time predictions. Note that the RF-GCN model has not been used for running time predictions, as zero observations are impossible here. Looking at Table IV, we can see a quite different picture than in Table III. This is not unexpected, as running time and dwell time have different characteristics. Here we can see that for almost all of the sites, the tree-based methods RF and XGBoost are superior, with XGBoost being the best model for Tampere and Graz, RF being the best for Madrid, and the performance being almost equal for Les Mureaux. Only for Linkoping, the ANN models outperform, with both the MLP and the GCN having comparable performance. For Tampere, the difference between the MLP and XGboost is only half a second, while the difference between Graz, Madrid, and Les Mureaux is around 2 seconds.
It can also be seen that the MLP achieves better results than the GCN for all sites. This is different than what has been observed with using GCNs for other traffic-related tasks, such as traffic flow, where in general, the spatial modeling process of GCNs improves performance [24]. A reason for this result could be the deployment in low-traffic zones or the regulatory speed limit for automated shuttles, leading to less variation in the running times.
@@
VI-A Assessment of dwell time predictions
@@
The previous section showed that the optimal choice between deep learning and tree-based methods for predicting dwell times is site-dependent, with no universally best model. In light of these findings, a deeper understanding is needed to determine the conditions that favor one approach over the other.
In Figure 3, we depict the dwell time distribution in Linköping. Linköping is unique in having the longest ongoing pilot and route, collecting nearly five times more observations than the other sites. The dwell time distribution here is distinctly multimodal. A mode at zero seconds indicates instances where a shuttle stop is skipped. The remaining modes suggest times when the shuttle stops and passengers are alighting or boarding. There is a variation of the dwell times between 20 and 40 seconds, indicative of a frequently used service. In Figure 3, we can observe a difference in the performance of the prediction models. Figure 2(a) illustrates that XGBoost struggles to categorize the majority of zero dwell times and overlooks the third mode at 35 seconds. In contrast, our hybrid RF-GCN (2(b)) demonstrates a better fit. The Random Forest classifier effectively classifies if a shuttle is stopping or skipping a stop, and the GCN model fits the multimodal dwell time distribution for cases when passengers are alighting and boarding. Remarkably, by incorporating spatial information, this model even differentiates between the two primary dwelling times of approximately 25 and 35 seconds.
On the other hand, the benefits of tree-based methods are evident in the case of the Les Mureaux pilot site (Figure 4). In Les Mureaux, we can notice the automated nature of the service with two main dwell times of 0 and 25 seconds for skipping and stopping, respectively. The data from this site suggests a default dwell time of 25 seconds that extends only when necessary. As prior research has shown [25], neural networks are biased towards overly smooth solutions. In contrast, decision trees, learning piece-wise constant functions, avoid this pitfall. This characteristic is evident in Figure 4, where we can observe the superiority of the Random Forest when fitting to a non-smooth target function.
Therefore, we can conclude that the relative advantage of using a GNN depends on the specific characteristics of the shuttle service. In cases with high data availability and a frequently-used service with a long route, the GNN seems superior. Yet, the random forest remains relevant in detecting the skipped stops. However, in scenarios where the dwell time is largely automated and does not significantly vary with passenger density, the prediction challenge simplifies to a classification task: Is the shuttle stopping or not? Here, the tree-based methods show a clear advantage.
@@
VI-B Evaluation on journey
So far, our focus has been on evaluating model performance over individual segments. However, in real-world service deployment, the cumulative AT prediction, formed by aggregating individual segment predictions, is more important. Hence, for each pilot site, we sample five journeys from the test set and assess the accumulated error along the journey starting from stop number 1. We illustrate the results for the three sites with the longest routes, Linköping, Les Mureaux, and Tampere, in Figures 5-7. Solid lines depict average accumulated errors, while the shaded regions represent the range between maximum positive and negative errors across the sampled journeys.
Linköping has the longest route of the pilot sites, with a total of 15 stops. In Figure 5, stop number 7 corresponds to a 15-minute ahead, and stop number 15 to a 35-minutes ahead prediction. The depicted models (RF-GCN, XGBoost, and the Mean regressor) overestimate the total travel time, but RF-GCN shows the least average error with a maximum of 30 seconds. In the 15-minutes ahead predictions at stop number 7, we can observe an error of a maximum of 60 seconds and 35 minutes ahead of around 100 seconds. XGBoost still shows a lower average error than the Mean regressor, yet, toward the end of the route, the accumulated errors display worse maximum values.
The interesting finding in the Tampere (Figure 7) site is that although XGboost outperforms the GCN on running time prediction, the accumulated predictions by the RF-GCN/GCN combination are more accurate. The average error is close to zero, and the maximum error stays below 100 seconds, while XGBoost is, on average, underestimating the running time by up to 50 seconds on average and 150 seconds in the worst case. These findings suggest that future work needs to rethink how they evaluate their models, as most existing work does not consider aggregation errors.
@@ VI-B Evaluation on journey
Since we are aggregating the predictions of the dwell and running time prediction models, the question arises of how the errors accumulate between those two models and which one is driving the difference in performance in Figures 5-7. In Table V, we calculate the average accumulated absolute error over the entire journey resulting from the dwell and the running time prediction, respectively. The performance of the mean regressor on Linköping and Tampere reveals that the error originating from the dwell time prediction has a larger share of the total travel time error than the one from the running time model. Similarly, we can identify the reason for the superior performance of the GCN in the accumulated travel times, as it decreases the error in dwell time prediction significantly while the error in running time prediction stays comparable. More specifically, if we compare the decrease of errors of the best model compared to the mean regressor, we observe a decrease of 35%, 26%, and 61% in dwell time error for the three pilot sites and a decrease of 17%, 15%, and 32% for the error originating from the running time predictions. With these results, we stress that dwell time prediction is an important factor for the AT prediction for autonomous shuttles. Especially when they are used in remote, low-traffic areas or are confined to regulatory speed limits, dwell time becomes the most influential factor.
TABLE V: Averaged Accumulated Absolute Error
@@
GCN 93 105 56 57 16 64
VI-C Discussion

### v2

@@ IV-A Data pre-processing
Finally, we enrich our data set with weather data at an hourly resolution and perform sine and cosine encodings of time-of-day and day-of-week features. For a correct model evaluation, we hold out a part (approximately 10 $\%$) of the data set depending on the date to avoid data leakage and test temporal generalisation. In Linköping, Tampere, Madrid, and Graz, our test set selects one month of data and one week in Les Mureaux and Carinthia.
@@
The data pre-processing step is of utmost importance to make an accurate distinction between dwell time and running time. Since our data originates from real-world pilot tests, we found careful data cleaning to be essential to obtain a data set that reflects regular operation as accurately as possible.
@@
TABLE III: Results for Running Time Predictions Depending on Lags (Results are (RMSE/MAE))
@@
MAPE 38.48 39.94 36.55 29.48 30.44 36.57 $\pm$ 0.29 37.03 $\pm$ 0.33
$R^{2}$ 0.08 0.37 0.43 0.51 0.52 $0.44\pm 0.00$ $0.48\pm 0.00$
Carinthia RMSE 31.52 26.53 26.50 27.16 26.07 $31.86\pm 0.14$ $29.68\pm 1.06$
MAE 20.74 16.94 17.16 17.60 16.78 $22.02\pm 0.09$ $19.84\pm 0.59$
MAPE 12.50 9.94 10.30 10.31 9.96 13.92 $\pm$ 0.76 13.67 $\pm$ 2.06
$R^{2}$ 0.81 0.87 0.87 0.86 0.87 $0.81\pm 0.02$ $0.83\pm 0.00$
V Results
In this section, we perform a series of experiments on the data sets presented in the previous section. First, we investigate whether it is beneficial to create historical lags based on all vehicles or per vehicle (since different vehicles may correspond to different models and manufacturers). Then we compare the performance of the different prediction models in running and dwell time prediction, and aggregate the predictions to calculate the final arrival time along observed journeys from the test data to evaluate compounding error and model bias. Lastly, we quantify the impact of deviations from the scheduled, automated service and provide some key insights for improved data collection during pilot deployments. As performance metrics, we report the Root Mean Square Error (RMSE), the Mean Absolute Error (MAE), the Mean Absolute Percentage Error (MAPE), as well as $R^{2}$:
[…]
where $y$ is the observed target values and $\hat{y}$ is the predictions for the test set with $N$ observations.
V-A Choice of lag feature
Prior to assessing the final prediction results, we need to determine whether lags are taken from the most recent observation of the individual vehicle or the whole fleet. Hence, in Table III, we compare the predictive performance of the Lag Model and XGBoost on running time depending on the lags. Pilot testing occasionally results in sparse data, making recent lags from the whole fleet potentially advantageous. If, on the other hand, the fleet is a mix of shuttles from different manufacturers, observations from different vehicles may not be as informative. As seen in Table III, the results are not consistent across sites. Les Mureaux and Tampere deploy the same type of vehicles, while the remaining sites test different manufacturers. As a consequence, we can see a clear improvement in Linköping, with the highest number of vehicles deployed, when providing lags from the respective vehicle. For the other pilot sites, the difference is less striking. In Les Mureaux, the vehicles are identical, which leads to a better lag model for the lags based on the whole fleet. Yet, XGBoost performs better at the individual vehicle level, indicating subtle differences in the data of the vehicles. Based on this analysis, we create the lags per vehicle in Linköping, Tampere, and Les Mureaux and for the whole fleet in Madrid.
Figure 3: Mean SHAP values for XGB running time prediction. Variables include: segm_id (segment ID), tod (time-of-day), temp (temperature), wspd (windspeed), dow (day-of-week), prcp (precipitation)
V-B Running time prediction
In evaluating prediction results on single segments, we first analyze the comparative performance of machine learning and deep learning models across pilot sites. In Table IV, we present the performance for running time predictions across models, excluding the RF-GCN model, as zero observations are impossible. For most pilot sites, the tree-based methods RF and XGBoost are superior, with XGBoost outperforming in Tampere and Graz, RF in Madrid, and the performance being equal in Les Mureaux. Only for Linköping, the ANN models outperform in MAE, with both the MLP and the GCN showing comparable performance.
Surprisingly, across sites, MLP consistently outperforms GCN with the exception of Carinthia. This is different from findings for other traffic-related tasks, such as traffic flow prediction, where, in general, the spatial modeling process of GCNs improves performance [Li2017DiffusionCR]. A reason for this result could be the deployment in low-speed zones or the regulatory speed limit for automated shuttles, leading to less variation in the running times.
However, we are not only interested in identifying the most competitive model; more importantly, we aim to understand the overall predictability of automated running times and how it varies across different pilot sites. Analyzing $R^{2}$ scores indicates that, for most cities, simple models like the Lag and Mean Regressors sufficiently explain the variability in running times. With the exception of Madrid, the improvement in prediction accuracy when using more sophisticated models compared to the Mean Regressor (or the Lag model in the case of Les Mureaux) is minimal. Low average speeds and automated driving behavior result in stable running times that simpler models can effectively capture. Madrid, however, stands out with its deployment at a bus depot, focusing on automatic depot management and parking, and therefore showing the lowest average running times (63s) and highest coefficient of variation (0.34). In this case, we can observe a clear benefit of using a prediction model, but even the MAPE and $R^{2}$ of the best model are significantly higher than for the other sites. The models on Graz show a lower $R^{2}$ score than in the remaining pilot sites, reflecting data quality issues. Graz has not only the lowest number of observations but, more importantly, the pilot is also scattered across 9 months, leading to data sparsity that leaves the lag feature outdated.
These findings are further supported by feature importance analysis with SHAP values [NIPS2017_8a20a862] for the XGBoost regressor in Figure 3. For all pilot sites except Madrid and Graz, the lag feature is the most influential. In Madrid, the poor predictability of the lag model forces the XGBoost model to rely more on positional data, while in Graz, segment encodings are prioritized due to outdated lags caused by data sparsity. Notably, in Les Mureaux, which operates on a private company site with the lowest average speed, the Lag Model achieves an impressive $R^{2}$ of 0.95, making the lag feature by far the dominant predictor for this location. This performance underscores the predictive strength of the lag feature under stable, automated conditions.
TABLE V: Results for Dwell Time Prediction
Lag Mean LR RF XGBoost MLP GCN RF-GCN
$R^{2}$ -0.43 0.12 0.24 0.28 0.35 $0.21\pm 0.00$ $0.31\pm 0.00$ $\textbf{0.36}\pm 0.00$
$R^{2}$ -0.55 0.07 0.1 0.15 0.16 $00\pm 0.01 0.0$$01\pm 0.01$ $0.14\pm 0.00$
MAE 5.41 8.51 5.85 3.30 3.46 $5.45\pm 0.11$ $5.23\pm 0.15$ $3.98\pm 0.13$
$R^{2}$ -0.13 -0.09 0.26 0.53 0.51 $0.26\pm 0.02$ $0.30\pm 0.03$ $0.47\pm 0.01$
Graz RMSE 15.84 10.88 11.27 10.52 10.66 $12.71\pm 0.24$ $2.92\pm 0.17 11$$1.21\pm 0.09$
$R^{2}$ -1.12 0.00 -0.07 0.07 0.04 $0.37\pm 0.05$ $41\pm-0.04 0.0$$06\pm 0.02$
$R^{2}$ -0.90 -0.37 -0.24 -0.18 0.05 $04\pm 0.02 0.0$$6\pm 0.00 0.0$$07\pm 0.01$
Carinthia RMSE 23.70 19.08 18.26 17.34 17.31 $18.90\pm 0.17$ $19.44\pm 0.40$ $18.12\pm 0.02$
[…]
$R^{2}$ -0.21 0.22 0.28 0.35 0.36 $0.23\pm 0.01$ $0.18\pm 0.03$ $0.29\pm 0.00$
Figure 4: Mean SHAP values for dwell time prediction with XGB.
V-C Dwell time prediction
In Table V, we analyze whether the results from running time predictions translate to dwell time predictions. Interestingly, there is no universal best model across all sites, showing quite a different picture from Table IV. This is not unexpected, as running time and dwell time have different characteristics. In Linköping, Tampere, and Carinthia, XGBoost outperforms in RMSE, while RF-GCN achieves the lowest MAE, with the difference in MAE being more significant than in RMSE. These results highlight the benefit of integrating spatial information for dwell time prediction in suburban residential areas, where these three pilots are situated. In Les Mureaux, operating employee shuttles, the situation is reversed. The tree-based models outperform the other models by a large margin. We perform a more detailed analysis of this difference in Section V-D, identifying automated dwell times as a contributing factor. Given the very limited data availability in Graz, it is no surprise that our tested models cannot outperform the Mean Regressor, as there is not enough data for the advanced models to learn the spatial and temporal correlations. In Madrid, the GCN demonstrates better performance than RF-GCN. We can attribute this to an observed degradation in the performance of the RF classifier in detecting zeros compared to the standalone GCN. This intriguing finding suggests that spatial information in this pilot, which does not focus solely on passenger transportation, not only helps in accurate dwell time prediction during stops but also in detecting stop skips.
Again, we are not only interested in the best-performing prediction models but also in understanding the general predictability of dwell times for automated shuttles. Unlike in running time prediction, both baseline models, Lag and Mean, yield low and sometimes even negative $R^{2}$ values across all pilot sites, highlighting dwell time prediction as the more challenging task. Given these results, it is surprising that most literature focuses on travel or running time prediction, while dwell time prediction has received far less attention. The fact that the lag feature is less informative for dwell time prediction than for running time prediction is also reflected in the feature importance analysis in Figure 4. Here, position stands out as the most influential feature by a large margin. As a result, the model must rely on a broader range of input features for accurate dwell time prediction, i.e., temperature and time-of-day features, which only had minimal impact on running time.
Evaluating the $R^{2}$ of the best-performing models reveals substantial differences across pilot sites, as expected given their different characteristics. In Graz and Madrid, the models barely achieve positive $R^{2}$ values. In Graz, this can largely be attributed to the limited data availability. Additionally, as the only pilot site using passenger cars, factors such as low-clearance boarding or alighting, along with passengers manually operating the doors, may cause dwell times to be more dependent on passenger characteristics (e.g., mobility) and the number of passengers (e.g., when the exit is limited to a single door). This situation seems to lead to harder-to-predict dwell times compared to conventional shuttle buses. These findings are supported by the results from Tampere, which show a lower $R^{2}$ score than Linköping or Carinthia, despite operating in areas with similar characteristics and having a comparable or higher number of observations but deploying a van instead of a bus. The poor results in Madrid are attributed to the focus on automated depot management, resulting in very high variability in dwell times as noted in the previous section. This variability appears to be driven more by the testing of operational procedures than by consistent passenger demand patterns, leading to greater unpredictability. Finally, in Les Mureaux, $R^{2}$ scores are significantly higher than in the other sites, which can be explained by an automated, default dwell time at this pilot site, a factor that will be explored in more detail in the next section. Overall, across all sites except Graz, using ML-based models significantly improves performance, demonstrating their value in dwell time predictions.
V-D Detailed analysis of dwell time predictions
@@
The previous section showed that the optimal choice between deep learning and tree-based methods for predicting dwell times is site-dependent, with no universally best model. In light of this, a deeper understanding is needed to determine the conditions that favor one approach over the other.
@@
In Figure 5, we depict the dwell time distribution in Linköping. Linköping is unique in having the longest ongoing pilot and route, collecting nearly five times more observations and serving more than double the number of passengers as the other sites. The dwell time distribution here is distinctly multimodal. A mode at zero seconds shows instances where a shuttle stop is skipped. The remaining modes suggest times when the shuttle stops and passengers are alighting or boarding. There is a variation of the dwell times between 20 and 40 seconds, indicative of a frequently used service. In Figure 5, we observe a difference in the performance of the prediction models. Figure 5(a) illustrates that XGBoost struggles to categorize the majority of zero dwell times and overlooks the third mode at 35 seconds. In contrast, our RF-GCN (5(b)) demonstrates a better fit. The RF classifier effectively classifies whether a shuttle is skipping a stop, and the GCN model fits the multimodal dwell time distribution for cases when passengers are alighting and boarding. Remarkably, by incorporating spatial information, this model even differentiates between the two primary dwell times of approximately 25 and 35 seconds.
On the other hand, the benefits of tree-based methods are evident in the case of Les Mureaux (Figure 6), where the shuttles are only deployed for employees on a private company site and thus shows the second lowest number of served passengers. Here, we can notice the automated nature of the service with two main dwell times of 0 and 25 seconds for skipping and stopping, respectively. The data from this site suggests a default dwell time of 25 seconds that extends only when necessary. As prior research has shown [Grinsztajn], neural networks are biased towards overly smooth solutions. In contrast, decision trees, learning piece-wise constant functions, avoid this pitfall. This characteristic is evident in Figure 6, where we can observe the superiority of the Random Forest when fitting to a non-smooth target function.
Therefore, we conclude that the relative advantage of using a GNN depends on the specific characteristics of the shuttle service. In cases with high data availability, the frequently-used service with a long route, and deployment in suburban, residential areas, the GNN seems superior. Yet, the random forest remains relevant in detecting the skipped stops. However, in scenarios where the dwell times are largely automated, do not significantly vary with passenger density, or service types with lower passenger counts (e.g., employee-only shuttles), the prediction challenge simplifies to a classification task: is the shuttle stopping or not? In these conditions, the tree-based methods show a clear advantage.
V-E Evaluation on journey
@@
So far, our focus has been on evaluating model performance over individual segments. However, in real-world service deployment, the cumulative AT prediction, formed by aggregating individual segment predictions, is more important. Hence, we sample five journeys from the test set and assess the accumulated error along the journey starting from stop number 1. We illustrate the results for the three sites with the longest routes, Linköping, Les Mureaux, and Tampere, in Figures 7-9. Solid lines depict average accumulated errors, while the shaded regions represent the range between maximum positive and negative errors across the sampled journeys. At each prediction instant, we freeze the features to whatever data is available while the shuttle is still at stop 1. Lag variables, therefore, reflect the most recent observations before the current trip begins; they are not updated for predictions further ahead. We evaluate the accumulated error across sample strips to evaluate systemic prediction bias and compounding errors.
Linköping has the longest route of the pilot sites, with a total of 15 stops. In Figure 7, stop number 7 corresponds to a 15-minute, and stop number 15 to a 35-minute ahead prediction. The depicted models (RF-GCN, XGBoost, Mean Regressor) overestimate the total travel time, but RF-GCN shows the least average error with a maximum of 30 seconds. In the 15-minute ahead predictions at stop number 7, we can observe an error of a maximum of 60 seconds and 35 minutes ahead of around 100 seconds. XGBoost still shows a lower average error than the Mean Regressor, yet, toward the end of the route, the accumulated errors display worse maximum values.
The interesting finding in the Tampere (Figure 9) site is that although XGboost outperforms the GCN on running time prediction, the accumulated predictions by RF-GCN/GCN are more accurate. The average error is close to zero, and the maximum error stays below 100 seconds, while XGBoost is underestimating the running time by up to 50 seconds on average and 150 seconds in the worst case. These findings suggest that future work needs to rethink model evaluation, as most existing works do not consider aggregation errors.
TABLE VI: Averaged Accumulated Absolute Error (s)
@@
GCN/GCN-RF 93 105 56 57 16 64
Since we are aggregating the predictions of the dwell and running time prediction models, the question arises of how the errors accumulate between those two models and which one is driving the difference in performance in Figures 7-9. In Table VI, we calculate the average accumulated absolute error over the entire journey resulting from the dwell and the running time prediction, respectively, which is calculated by $\frac{1}{|J|}\;\sum_{j=1}^{J}\;\sum_{k=1}^{N_{j}}\Bigl|{\hat{s}_{k}}\;-\;s_{k}\Bigr|$, where $J$ is the set of Journeys and $N_{j}$ is the set of segments or stops in journey $j$, $\hat{s}_{k}$ the predicted time for segment $k$, and $s_{k}$ the observed time for segment $k$. The performance of the Mean Regressor on Linköping and Tampere reveals that the error originating from the dwell time prediction has a larger share of the total travel time error than the one from the running time model. Similarly, we can identify the reason for the superior performance of the GCN in the accumulated travel times, as it decreases the error in dwell time prediction significantly while also further decreasing the error in running time prediction.
Interestingly, while evaluating individual segments in Section V-B did not show a clear advantage for using a prediction model for running time, the accumulated errors reveal a notable improvement when the errors are accumulated over the entire journey. More specifically, if we compare the decrease of errors of the best model to the Mean Regressor, we observe a decrease of 35%, 26%, and 61% in dwell time error for the three pilot sites and a decrease of 17%, 15%, and 32% for the error originating from the running time predictions. With these results, we stress that dwell time prediction is an important factor for the AT prediction for automated shuttles. Especially when they are used in remote, low-speed areas or are confined to regulatory speed limits, dwell time becomes the most influential factor.
Figure 10: Stop location in the Carinthia pilot site (blue) and locations with opened door (red)
V-F Deviation from scheduled, automated service
Some pilot sites have reported allowing safety drivers to make unscheduled stops to let passengers board in an effort to increase public acceptance. This manual on-demand service, being unplanned, is not recorded in the datasets. For one of the pilot sites, Carinthia, we, however, do have access to door status information, which allows us to track stops outside the scheduled service. Figure 10 shows the scheduled stops alongside all door-opening occurrences, highlighting the challenges of working with real-world data. Our arrival time prediction approach is designed for scheduled, automated shuttle services, and these unplanned stops introduce variability in running times that is not captured by our features, such as time of day or weather conditions, but rather depends on the decisions of the safety driver. Specifically, we observe that unplanned stops occurred during $20\%$ of the running time segments, adding noise to the overall running time calculation and increasing complexity for prediction models. To provide further insight, Figure 11 shows the reduction in variability (i.e., in standard deviation) by filtering out the segments with stops en-route. For longer segments, such as 2_3 or 6_7, unplanned stops account for more than $30\%$ of the running time variability rather than the actual driving time.
This variability impacts prediction model performance. By evaluating on the filtered dataset, we observe a decrease in MAE of $1.8\%$ for the Mean Regressor, $4.5\%$ for XGBoost, and $9.2\%$ for the Random Forest regressor. These results show that machine learning methods are more susceptible to the additional noise introduced by unplanned stops, highlighting the importance of filtering or accounting for such variability to improve prediction accuracy.
Figure 11: Comparison of standard deviation in running times when filtering out on-demand service
V-G Key insights from working with data from real-world automated shuttle deployment